OpenStudy (anonymous):

how can i make -1+sqrt(3i) into trig form

OpenStudy (jhannybean):

@wio let's solve this!!

OpenStudy (jhannybean):

Never tried these types of problems before, I'm interested. I don't know where to begin though, lol.

OpenStudy (anonymous):

It is\[ -1+\sqrt 3 i \]Or is it: \[ -1+\sqrt{3i} \]

OpenStudy (anonymous):

I believe the \(i\) should not be in the square root, but if it is we can still do this.

OpenStudy (jhannybean):

Shall we try it both ways?

OpenStudy (anonymous):

Okay, we can do that.

OpenStudy (anonymous):

So the trig form would is like: \[ a+bi = re^{i\theta} = r[\cos(\theta) + i\sin(\theta)] \]

OpenStudy (jhannybean):

http://openstudy.com/study#/updates/530bfd53e4b03fec1e318cfc was looking at previous examples.

OpenStudy (anonymous):

Basically, \((a,b)\) represent the Cartesian coordinates, while the polar coordinates are \((r,\theta)\).

OpenStudy (jhannybean):

Oh i see.

OpenStudy (anonymous):

So: \[ r = \sqrt{a^2+b^2} \\ \theta =\tan^{-1}\left(\frac ab\right) \]

OpenStudy (anonymous):

It really is that simple.

OpenStudy (jhannybean):

I thought it'd be a little bit more complex...haha.

OpenStudy (anonymous):

One thing to notice is that \(a^2+b^2 = z\bar z\) where \(z=a+bi\) and \(\bar z=a-bi\) (called the conjugate).

OpenStudy (anonymous):

We call \(r=\sqrt{z\bar z}\) the modulus or \(z\), for some reason, and we call \(\theta\) the argument.

OpenStudy (jhannybean):

Oh okay. Makes sense.

OpenStudy (anonymous):

so \[ r = \sqrt{1^2+\sqrt{3}^2} = \sqrt{1+3} = 2 \]

OpenStudy (anonymous):

Whoops, I messed up earlier. I should have said: \[ \theta = \tan^{-1}\left(\frac ba\right) \]

OpenStudy (jhannybean):

b/a?... Oh yeah, similar to y/x

OpenStudy (anonymous):

In fact, it would be easier to just say \[ z=x+yi \]and use \((x,y)\). By convention \((a,b)\) is more popular.

OpenStudy (anonymous):

I don't like complex analysis, because we should just be learning about 2D vectors instead.

OpenStudy (jhannybean):

Hmm.

OpenStudy (anonymous):

The only thing about complex analysis that is interesting is complex multiplication. There isn't any vector equivalent.

OpenStudy (anonymous):

Well, maybe a couple other things as well. I never fully did complex analysis.

OpenStudy (jhannybean):

I didn't know that's what this was xD

OpenStudy (anonymous):

Anyway \[ \theta = \tan^{-1}\left(\frac{\sqrt3}{-1}\right)=-\frac{\pi}{3} \left(=-60^\circ\right) \]

OpenStudy (anonymous):

\[ -1+\sqrt 3i = 2\left[\cos\left(-\frac\pi 3\right)+i\sin\left(-\frac\pi 3\right)\right] \]

OpenStudy (jhannybean):

Wait, why is \(a=1\)? Wouldn't \(a=1\)?

OpenStudy (anonymous):

\[ r = \sqrt{(-1)^2+(\sqrt{3})^2} = \sqrt{1+3} = 2 \]

OpenStudy (jhannybean):

Oh that's right. Okay.

OpenStudy (anonymous):

In the case of: \[ -1+\sqrt{3i} \]You basically have to find the \(\sqrt{i}\), which actually has two solutions.

OpenStudy (mendicant_bias):

Heyoooo, Euler's Formula, I *do* recognize at least some of this, lol.

OpenStudy (anonymous):

\[ i = (x+yi)^2=(x^2-y^2)+(2xy)i \]

OpenStudy (anonymous):

This means\[ x^2-y^2=0,\quad 2xy = 1 \]

OpenStudy (anonymous):

since we can say \(i = (0) + (1)i\)

OpenStudy (jhannybean):

I'm little stuck! not understanding where the 0 and 1 are coming from...>.<

OpenStudy (anonymous):

Well, basically \(i\) is a complex number of the form \(x+yi\) where \((x,y) = (0,1)\).

OpenStudy (jhannybean):

Oh okay. I thought you derived it from something else. Nvm!

OpenStudy (anonymous):

\[ x^2-y^2=0\implies x^2=y^2 \implies |x|=|y| \implies x=y,x=-y \]Plug it into the other equation an: \[ 2yy=1,2(-y)y = 1 \]

OpenStudy (jhannybean):

Oh cool

OpenStudy (anonymous):

We can't let \(x\) or \(y\) be imaginary numbers, so the only equation that matters is: \[ 2y^2=1 \implies y=\pm \sqrt{\frac{1}{2}} \]

OpenStudy (jhannybean):

Why can we not let them be imaginary?

OpenStudy (anonymous):

Because they come from: \[ x+yi\], which is already a complex number, with real coefficients.

OpenStudy (jhannybean):

Ohh. I see that connection.

OpenStudy (anonymous):

Okay, so let's test it out real quick: \[ \left(\frac 1{\sqrt 2} + \frac 1{\sqrt{2}}i \right)^2 \]

OpenStudy (anonymous):

\[ =\frac 12 +\frac 12i+\frac 12i -\frac 12 = \left(\frac 12 -\frac 12\right) +\left(\frac 12+\frac 12\right)i = 0+1i = i \]

OpenStudy (anonymous):

If that is a solution, then so is: \[ -\frac 1{\sqrt2}-\frac{1}{\sqrt 2}i \]

OpenStudy (jhannybean):

nvm that would result in something completely different.

OpenStudy (anonymous):

No, if \(z^2=i\) then \((-z)^2=i\). This works for complex numbers as well.

OpenStudy (jhannybean):

Ahh, interesting. O_o

OpenStudy (anonymous):

We ended up with\[ x=y\\ 2y^2=1 \]So we know \(x\) and \(y\) must have the same sign and absolute value.

OpenStudy (jhannybean):

I see. Got it.

OpenStudy (anonymous):

Another way to do this, with less algebra, is to use trig coordinates.

OpenStudy (anonymous):

\[ i=(0)+i(1)=1[\cos(\pi/2)+i\sin(\pi/2)]=1e^{i\pi/2} = e^{i\pi/2} \]

OpenStudy (anonymous):

Take the square root, and we get: \[ \pm\sqrt{e^{i\pi/2}} = \pm e^{i\pi/4} = \pm[\cos(\pi/4)+i\sin(\pi/4)] = \pm\left[\frac 1{\sqrt2}+\frac 1{\sqrt2}i\right] \]

OpenStudy (anonymous):

The only reason all of this works is due to the fact that: \[ [\cos(\theta)+i\sin(\theta)]^n = \cos(n\theta)+i\sin(n\theta) \]

OpenStudy (anonymous):

And this formula works due to than double angle formula: \[ \cos(2\theta) = \cos^2(\theta)-\sin^2(\theta)\\ \sin(2\theta) = 2\cos(\theta)\sin(\theta) \]

OpenStudy (jhannybean):

Ah, I don't think i'm comfortable with the function of \(i\) as of yet. I have to keep referencing back to your post describing what it is. \[i = (x+yi)^2=(x^2-y^2)+(2xy)i \iff x^2-y^2=0,\quad 2xy = 1\]

OpenStudy (jhannybean):

I understand how it equates to 0 and 1, but now i'm confused with how you got to \[=1[\cos(\pi/2)+i\sin(\pi/2)]=1e^{i\pi/2} = e^{i\pi/2}\] That part.

OpenStudy (anonymous):

Well, I wanted:\[ (0)+i(1) = \cos(\theta) + i\sin(\theta) \]where \[ \cos(\theta) =0,\quad \sin(\theta) =1 \]

OpenStudy (anonymous):

The only places where \(\cos(\theta)=0\) (in the first revolution \(\theta \in [0,2\pi)\)), is \(\theta=\pi/2\) and \(\theta=3\pi/2\).

OpenStudy (jhannybean):

Ok, i get that.

OpenStudy (anonymous):

That gives \(\sin(\pi/2)=1\) and \(\sin(3\pi/2)=-1\)... so we say \(\theta=\pi/2\).

OpenStudy (anonymous):

Then \[ (0)+i(1) = \cos(\pi/2)+i\sin(\pi/2). \]

OpenStudy (jhannybean):

Understood.

OpenStudy (anonymous):

I basically just converted \(i\) to its trig coordinates. Since you have sort of learned it a moment ago, I figured you would get it. But you're not completely familiar yet so I was going too fast.

OpenStudy (jhannybean):

Yeah a little bit.

OpenStudy (jhannybean):

But after that explanation I found the connection.

OpenStudy (anonymous):

Actually, for \(i\) my method for finding \(\theta\) breaks down a bit since you are using: \[ \theta = \tan^{-1}\left(\frac 10\right) \]

OpenStudy (jhannybean):

I see.

OpenStudy (anonymous):

You just have to realize that when \[ \tan(\theta) = \pm \infty \]Then it gives you \[ \theta = \pm \pi/2 + 2\pi n \]

OpenStudy (jhannybean):

Well yeah.  I just had to visualize it.